OpenStudy (anonymous):

Suppose you are tossing 3 fair six-sided dice 1,080 times. How many times will you expect a total score of ‘6’ to occur out of the 1,080 tosses?

OpenStudy (anonymous):

A total score of '6' occurs with probability 9/216 = 1/24. Out of 1080 tosses, we can expect a total score of '6' for 1080*1/24 = 45 times.